OpenStudy (anonymous):

A remainder of 0 in the process of synthetic division tells you that you have found a root, or a factor, of the polynomial. True or false?

OpenStudy (anonymous):

I said false

OpenStudy (cruffo):

It's actually true.

OpenStudy (anonymous):

Ohh. how?

myininaya (myininaya):

true!

OpenStudy (cruffo):

have you done any problems using synthetic division?

OpenStudy (anonymous):

yes. a while back

myininaya (myininaya):

f(x)=x^2-4 2| 1 0 -4 | 2 4 _______ 1 2 | 0 since we have 0 in that remainder spot then x=2 is a zero which also implies x-2 is a factor since x-2=0 actually the factored form is f(x)=x^2-4=(x-2)(x+2)